# Benchmarking Unsupervised Online IDS for Masquerade Attacks in CAN

### v2

@@ V-A Evaluation Metrics Contrast
Figs. 4, 5, 6, 7 show the detection performance of each detection algorithm (i.e., Matrix Correlation Distribution, Matrix Correlation Correlation, Ganesan17, and Moriano22, respectively). Each cell in the heatmaps depicts a single AUC-ROC value for a specific combination of $\omega$ and $\delta$. Note that attack types are displayed horizontally. As attack types have different number of associated captures (see Table I), we compute the average in each attack category and report a single heatmap summarizing the performance in a specific category. As the standard deviation from this aggregation is negligible with respect to the average, we do not present associated figures here. The results are rounded at two decimal places. Specifically, we vary the window length (i.e., $\omega$ from 50 to 400 samples in increments of 50 samples) and the offset (i.e., $\delta$ from 10 to $\omega$ in increments of 10 samples).
@@
Masquerade attacks in the ROAD dataset are sustained and stealthy, lasting about 17–72 seconds (see Table I). We tuned the sliding window parameters ($\omega$ and $\delta$) to match attack duration while keeping computation manageable. For example, a window size of 400 samples gives enough temporal context to detect slow-developing anomalies. An offset of 380 samples allows overlap between windows, which helps avoid missed detections at window edges. This offset does not delay response but balances detection coverage and computational cost. Smaller offsets (e.g., 100) increase load, while larger ones (e.g., 400) may miss attacks. These parameters match the timing of real attacks. Note that masquerade attacks do not show up in single CAN frames. They imitate normal traffic in short bursts but become apparent as gradual deviations. As shown by Jedh et al. [66], detecting these attacks requires looking across multiple frames, not just isolated ones.
@@
Fig. 7 shows results for the Moriano22 Method. We observe that the maximum AUC-ROC among attack types is achieved with the max_engine_ attack achieving an AUC-ROC of $1.00$ for a few combinarions of $\omega$ and $\delta$. Additionally, for this attack, we found a region where AUC-ROC rose $0.70$ that is contained at lower values of $\omega$ and $\delta$, i.e., $100\leq\omega\leq 400$ and $10\leq\delta\leq 160$. For the light_off_attack, we also observe a performance region with values of AUC-ROC above $0.70$, generally contained at lower values of $\omega$ and $\delta$, i.e., $\omega\leq 200$ and $\delta\leq 200$. This method also displays AUC-ROC values above $0.60$ for the light_on_attack generally contained at the entire domain of values of $\omega$ and $\delta$. In addition, we observe that the evaluation metrics for the remaining attacks (i.e., correlated_attack and max_speedometer_attack) are not as good as in the previous attacks barely reaching above $0.50$.
@@
V-B Evaluation Metrics Summary
Table II summarizes results extracted from the heatmaps. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the performance of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) and maximum ($\max$). We also show the average and standard deviations from the maximum values for each attack detection methods and attack types in the ROAD dataset. They are displayed in the last row and column in the table.
@@ V-B Evaluation Metrics Summary
TABLE II: Summary of evaluation metrics based on the AUC-ROC heatmaps. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the maximum values obtained by each attack and detection method denoted by $\overline{max}$ and $\sigma_{max}$.
@@
Finally, we performed grid search over predefined hyperparameter ranges for Moriano22. Specifically, we varied $r\in\{-5,-3,\ldots,3\}$ and $\alpha\in\{0.1,0.2,\dots,0.8\}$. These ranges cover a variety of standard configurations. To ensure a thorough evaluation, we applied this procedure to the specific combination of $\omega$ and $\delta$, where Moriano22 performed the worst among all attack scenarios. This allowed us to analyze the worst-case scenario and assess the performance improvements achievable by tuning hyperparameters. We visualized the sensitivity of AUC-ROC scores to these parameters using heatmaps. For each attack category, we computed the average performance and summarized it in a single heatmap. This approach helps interpret the effect of hyperparameter tuning on the method’s performance. Figure 9 illustrates how AUC-ROC scores improve with optimized combinations of $r$ and $\alpha$. Notably, scores increased significantly from their minimum values (see Table 2 column Moriano22) to their maximum values shown in the heatmaps, except for the light_off_attack category. In addition, Table III compares the AUC-ROC values of the Moriano22 method before and after hyperparameter optimization. This analysis confirms the robustness of the proposed method under optimized settings while highlighting areas for further optimization.
@@
TABLE III: Comparison of AUC-ROC values for the Moriano22 method before and after hyperparameter optimization. The “before” column shows values from Table 2, where the method produces the lowest result. The “after” column presents values obtained using the window length and offset combination from the “before” column. The corresponding $r$ and $\alpha$ combinations that achieve the maximum values are shown in parentheses. Attack types are listed in the rows, while results before and after optimization are in the columns. The last column reports the absolute percentage change.
@@
V-C Performance Metrics Summary
Table IV summarizes TTW results based on heatmaps that we do not display here. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the TTW of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). In addition, we also show the average and standard deviations from the minimum values for each attack detection methods and attack types in the ROAD dataset. They are displayed as the last row and column in the table.
@@ V-C Performance Metrics Summary
TABLE IV: Summary of performance metrics based on TTW in miliseconds. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the minimum values obtained by each attack and detection method denoted by $\overline{min}$ and $\sigma_{min}$.
@@
VI Discussion
In this work, we have benchmarked non-DL-based unsupervised online IDS for detecting masquerade attacks in CAN data. Our results indicate that the bechmarked algorithms are not equally effective at detecting every attack type as it has been shown before in previous studies in the offline setting [39, 18, 40]. We attributed this to the foundational principle used for each of the detection algorithms and the magnitude of the perturbation that is incurred when introducing each of the attacks. Overall, Moriano22 produces the best detection results across attack types. Among attacks, max_speedomer_attack and light_on_attack are the overall hardest to detect attack type.
@@ VI Discussion
Although Moriano22 outperformed the other methods, achieving the best detection results across attack types, it showed lower average AUC-ROC values for the correlated_attack (i.e., $\mu$ = 0.35). This result likely stems from the windowing approach used in the testing, which limits capturing longer-term effects of masquerade attacks. These longer-term effects may hinder the formation of consistent cluster hierarchies and could be further explored with other time series prequential evaluation methods [47]. However, Moriano22 outperformed the other methods on the remaining attacks—max_engine_attack, max_speedometer_attack, light_off_attack, and light_on_attack. This indicates greater robustness in detecting masquerade attacks, particularly in the presence of subtle data manipulations such as those affecting the engine coolant, speedometer, and reverse light signals. Given the variability in performance across attacks, we expect that combining the benchmarked algorithms through ensemble learning could improve overall robustness. An ensemble could exploit the strengths of individual methods and compensate for their weaknesses, especially when detecting subtle or sustained attacks.
@@
While previous research has focused on detecting masquerade attacks in an offline setting, our results demonstrate the importance of considering how sliding windows are processed in an online setting with respect to window length and offset. Specifically, our results provide new insights on the effect of choosing carefully these two data partitioning parameters. To make benchmarked methods more effective to work in a real-world setting, online algorithms should strive to perform well when processing small batches of data (i.e., lower $\omega$ values) and low latency between window transitions (i.e., lower $\delta$ values). We quantify these relationships for every method on each of the attack types.
@@
Our study provides insights on the effectiveness and performance of online detection algorithms for detecting masquerade attacks from the ROAD dataset. Having noted the strengths of our benchmark framework, we are also aware of the following limitations of our proposed work.
@@
Lack of Validation in a Real Environment: Our proposed framework partitions logs of already collected CAN captures to simulate an streaming environment of CAN data using windows. We did not empirically validated our results on a moving vehicle including the necessary hardware and performance tuning involved to do so.
@@
Limited Number of Benchmarked Algorithms: We bechmarked four unsupervised online detection algorithms. As we strive for computationally efficient approaches and test their efficacy on a widely known attack dataset, we did not include unsupervised DL-based approaches (e.g., [19, 16]) in the comparison.
Focus on a Single Testbed Dataset: We used the ROAD dataset with a limited number of masquerade attacks to benchmark online detection algorithms. We did not consider newer datasets with a higher number of masquerade attacks generated on modern automobiles [68].
@@
VII Conclusion
This paper benchmarks non-DL-based unsupervised online IDS for detecting masquerade attacks in CAN. We focus on four detection algorithms (i.e., Matrix Correlation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]) which foundational mechanism is based on capturing significant deviations from expected pairwise correlations over a stream of multivariate time series. The inclusion of these baseline algorithms along with their open source implementation allow direct comparison with forthcoming online algorithms in this research space. The presented benchmark differs from other benchmarks in prior works in that it focuses on the online operation of detection algorithms (as opposed to offline operation mode) aimed at detecting masquerade attacks—the stealthiest category of attacks in the CAN bus. Our benchmark is based on the concept of time windows that control the partition of a stream of time series. Specifically, we use sliding windows, with respect to the number of observations, to analyze variations in the AUC-ROC based on different combinations of window length and offset. We report conditions that makes benchmarked algorithms more successful than others for certain attack types as well a summary of their overall performance on different attacks. We show that overall, Moriano22 performs the best at detecting the different attack types in the ROAD dataset. Our analysis reveals that the max_speedometer_attack is the more elusive to detect among detection algorithms.

### v3

@@ 5.1 Evaluation Metrics Contrast
Figs. 4, 5, 6, 7 show the detection performance of each detection algorithm (i.e., Matrix Correlation Distribution, Matrix Correlation Correlation, Ganesan17, and Moriano22, respectively). Each cell in the heatmaps depicts a single AUC-ROC value for a specific combination of $\omega$ and $\delta$. Note that attack types are displayed horizontally. As attack types have different number of associated captures (see Table 1), we compute the average in each attack category and report a single heatmap summarizing the performance in a specific category. As the standard deviation from this aggregation is negligible with respect to the average, we do not present associated figures here. The results are rounded at two decimal places. Specifically, we vary the window length (i.e., $\omega$ from 50 to 400 samples in increments of 50 samples) and the offset (i.e., $\delta$ from 10 to $\omega$ in increments of 10 samples).
@@
Masquerade attacks in the ROAD dataset are sustained and stealthy, lasting about 17–72 seconds (see Table 1). We tuned the sliding window parameters ($\omega$ and $\delta$) to match attack duration while keeping computation manageable. For example, a window size of 400 samples gives enough temporal context to detect slow-developing anomalies. An offset of 380 samples allows overlap between windows, which helps avoid missed detections at window edges. This offset does not delay response but balances detection coverage and computational cost. Smaller offsets (e.g., 100) increase load, while larger ones (e.g., 400) may miss attacks. These parameters match the timing of real attacks. Note that masquerade attacks do not show up in single CAN frames. They imitate normal traffic in short bursts but become apparent as gradual deviations. As shown by Jedh et al. [71], detecting these attacks requires looking across multiple frames, not just isolated ones.
@@
Fig. 7 shows results for the Moriano22 Method. We observe that the maximum AUC-ROC among attack types is achieved with the max_engine_ attack achieving an AUC-ROC of $1.00$ for a few combinations of $\omega$ and $\delta$. Additionally, for this attack, we found a region where AUC-ROC rose $0.70$ that is contained at lower values of $\omega$ and $\delta$, i.e., $100\leq\omega\leq 400$ and $10\leq\delta\leq 160$. For the light_off_attack, we also observe a performance region with values of AUC-ROC above $0.70$, generally contained at lower values of $\omega$ and $\delta$, i.e., $\omega\leq 200$ and $\delta\leq 200$. This method also displays AUC-ROC values above $0.60$ for the light_on_attack generally contained at the entire domain of values of $\omega$ and $\delta$. In addition, we observe that the evaluation metrics for the remaining attacks (i.e., correlated_attack and max_speedometer_attack) are not as good as in the previous attacks barely reaching above $0.50$.
@@
To assist interpretation, we also provide Table 2, which report average trends across configurations.
5.2 Evaluation Metrics Summary
Table 2 summarizes results extracted from the heatmaps. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the performance of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) and maximum ($\max$). We also show the average and standard deviations from the maximum values for each attack detection methods and attack types in the ROAD dataset. They are displayed in the last row and column in the table.
@@ 5.2 Evaluation Metrics Summary
Table 2: Summary of evaluation metrics based on the AUC-ROC heatmaps. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the maximum values obtained by each attack and detection method denoted by $\overline{max}$ and $\sigma_{max}$.
@@
Finally, we performed grid search over predefined hyperparameter ranges for Moriano22. Specifically, we varied $r\in\{-5,-3,\ldots,3\}$ and $\alpha\in\{0.1,0.2,\dots,0.8\}$. These ranges cover a variety of standard configurations. To ensure a thorough evaluation, we applied this procedure to the specific combination of $\omega$ and $\delta$, where Moriano22 performed the worst among all attack scenarios. This allowed us to analyze the worst-case scenario and assess the performance improvements achievable by tuning hyperparameters. We visualized the sensitivity of AUC-ROC scores to these parameters using heatmaps. For each attack category, we computed the average performance and summarized it in a single heatmap. This approach helps interpret the effect of hyperparameter tuning on the method’s performance. Figure 9 illustrates how AUC-ROC scores improve with optimized combinations of $r$ and $\alpha$. Notably, scores increased significantly from their minimum values (see Table 2 column Moriano22) to their maximum values shown in the heatmaps, except for the light_off_attack category. In addition, Table 3 compares the AUC-ROC values of the Moriano22 method before and after hyperparameter optimization. This analysis confirms the robustness of the proposed method under optimized settings while highlighting areas for further optimization.
@@
Table 3: Comparison of AUC-ROC values for the Moriano22 method before and after hyperparameter optimization. The “before” column shows values from Table 2, where the method produces the lowest result. The “after” column presents values obtained using the window length and offset combination from the “before” column. The corresponding $r$ and $\alpha$ combinations that achieve the maximum values are shown in parentheses. Attack types are listed in the rows, while results before and after optimization are in the columns. The last column reports the absolute percentage change.
@@
5.3 Performance Metrics Summary
Table 4 summarizes TTW results based on heatmaps that we do not display here. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the TTW of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). In addition, we also show the average and standard deviations from the minimum values for each attack detection methods and attack types in the ROAD dataset. They are displayed as the last row and column in the table.
@@ 5.3 Performance Metrics Summary
Table 4: Summary of performance metrics based on TTW in miliseconds. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the minimum values obtained by each attack and detection method denoted by $\overline{min}$ and $\sigma_{min}$.
@@
6 Discussion
In this work, we have comparative evaluated non-DL-based unsupervised online IDS for detecting masquerade attacks in CAN data. Our results indicate that the evaluated algorithms are not equally effective at detecting every attack type as it has been shown before in previous studies in the offline setting [41, 18, 42]. We attributed this to the foundational principle used for each of the detection algorithms and the magnitude of the perturbation that is incurred when introducing each of the attacks. Overall, Moriano22 produces the best detection results across attack types. Among attacks, max_speedomer_attack and light_on_attack are the overall hardest to detect attack type.
@@ 6 Discussion
Although Moriano22 outperformed the other methods, achieving the best detection results across attack types, it showed lower average AUC-ROC values for the correlated_attack (i.e., $\mu$ = 0.35). This result likely stems from the windowing approach used in the testing, which limits capturing longer-term effects of masquerade attacks. These longer-term effects may hinder the formation of consistent cluster hierarchies and could be further explored with other time series prequential evaluation methods [52]. However, Moriano22 outperformed the other methods on the remaining attacks, i.e.,max_engine_attack, max_speedometer_attack, light_off_attack, and light_on_attack. This indicates greater robustness in detecting masquerade attacks, particularly in the presence of subtle data manipulations such as those affecting the engine coolant, speedometer, and reverse light signals. Given the variability in performance across attacks, we expect that combining the comparative evaluationed algorithms through ensemble learning could improve overall robustness. An ensemble could exploit the strengths of individual methods and compensate for their weaknesses, especially when detecting subtle or sustained attacks.
@@
While previous research has focused on detecting masquerade attacks in an offline setting, our results demonstrate the importance of considering how sliding windows are processed in an online setting with respect to window length and offset. Specifically, our results provide new insights on the effect of choosing carefully these two data partitioning parameters. To make comparative evaluationed methods more effective to work in a real-world setting, online algorithms should strive to perform well when processing small batches of data (i.e., lower $\omega$ values) and low latency between window transitions (i.e., lower $\delta$ values). We quantify these relationships for every method on each of the attack types.
While we recognize the value of deep learning-based IDSs, especially those validated on preselected signals or simplified environments, our comparative evaluation deliberately focuses on unsupervised non-DL methods that can operate on high-dimensional, unlabeled signal-level data without requiring prior semantic knowledge. Given the variability and scale of the ROAD dataset, along with the challenges in training, tuning, and deploying DL models in production vehicular environments, we restrict our analysis to more transparent and resource-efficient methods. Our released code provides an open pathway for future researchers to incorporate and evaluate DL-based methods under identical streaming configurations.
We acknowledge several limitations of our study, which we elaborate in the following section.
7 Threats to Validity
@@
Our study provides insights on the effectiveness and performance of online detection algorithms for detecting masquerade attacks from the ROAD dataset. Having noted the strengths of our comparative evaluation framework, we are also aware of the following limitations of our proposed work.
@@
Lack of Validation in a Real Environment: Our proposed framework partitions logs of already collected CAN captures to simulate an streaming environment of CAN data using windows. We did not empirically validated our results on a moving vehicle including the necessary hardware and performance tuning involved to do so. We acknowledge that true in-vehicle IDS deployment would require additional validation on automotive-grade microcontrollers. Our goal here is to evaluate algorithmic performance under realistic timing constraints using resource-constrained proxies, which is a common practice in CAN IDS literature.
@@
Limited Number of Comparative Evaluationed Algorithms: We evaluated four unsupervised online detection algorithms. As we strive for computationally efficient approaches and test their efficacy on a widely known attack dataset, we did not include unsupervised DL-based approaches (e.g., [19, 16]) in the comparison.
Focus on a Single Testbed Dataset: We used the ROAD dataset with a limited number of masquerade attacks to comparative evaluation online detection algorithms. We did not consider newer datasets with a higher number of masquerade attacks generated on modern automobiles [73]. While we focus on the ROAD dataset due to its broad signal coverage (hundreds of IDs and signals) and real driving instrumentation, we acknowledge that other datasets like CrySyS (Gazdag et al. [74]) offer complementary perspectives. CrySyS introduces signal modification on up to two signals in both physical and simulated environments, and is a valuable resource for future cross-dataset evaluation. Our modular comparative evaluation infrastructure and public code allow straightforward extension to include CrySyS or similar datasets. We consider this a promising avenue for subsequent work.
@@
8 Conclusion
This paper presents a comparative evaluation of non-DL-based unsupervised online IDS for detecting masquerade attacks in CAN. We focus on four detection algorithms (i.e., Matrix Correlation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]) which foundational mechanism is based on capturing significant deviations from expected pairwise correlations over a stream of multivariate time series. The inclusion of these baseline algorithms along with their open source implementation allow direct comparison with forthcoming online algorithms in this research space. The presented evaluation differs from other evaluations in prior works in that it focuses on the online operation of detection algorithms (as opposed to offline operation mode) aimed at detecting masquerade attacks, i.e., the stealthiest category of attacks in the CAN bus. Our evaluation is based on the concept of time windows that control the partition of a stream of time series. Specifically, we use sliding windows, with respect to the number of observations, to analyze variations in the AUC-ROC based on different combinations of window length and offset. We report conditions that makes evaluated algorithms more successful than others for certain attack types as well a summary of their overall performance on different attacks. We show that overall, Moriano22 performs the best at detecting the different attack types in the ROAD dataset. Our analysis reveals that the max_speedometer_attack is the more elusive to detect among detection algorithms.